# Controlling discrete semilinear wave equations toward flocks

### v1

@@ Abstract.
In this work, we initiate the research on controlling nonlinear waves propagating on lattices from a completely new perspective. We consider nonlinear waves on a lattice as a system of interacting particles and study their collective flocking behavior. By designing suitable feedback controls, we show that any admissible flock can be reached within a finite amount of time. Finally, we highlight the connection between our flocking problem and a minimal-time problem in the framework of nonlinear Hamilton–Jacobi equations and optimal control theory.
@@
The work of SS is partially supported by … The work of HVT is partially supported by NSF grant DMS-2348305. The work of MBT is partially supported by NSF Grants CAREER DMS-2303146, DMS-2204795, DMS-2306379.
@@ 1.1. Brief overview
In recent years, many efforts have been made to understand atomic vibrations in anharmonic chains and lattices. Atomic vibrations in anharmonic chains (one-dimensional systems) and lattices (multi-dimensional systems) can be described via discrete wave-type equations [5, 9, 11].
Control theory, which has been extensively studied since the last 60 years, has provided insight into several fields, including finance, medicine, mechanics, engineering, and climatology. Over several years, control theory has significantly shaped the study of wave equations [14].
The goal of our paper is to initiate the research on controlling atomic vibrations in anharmonic chains and lattices via the associated nonlinear discrete wave equations. In particular, nonlinear discrete wave equations can be applied to lattices in solid mechanics such as Peierls’ lattice [9] or to vibrating strings of oscillators like the Fermi-Pasta-UIam-Tsingou lattice [5]. We note that these lattices have a fixed grid size. It is well-known that wave vibrations on lattices can exhibit complicated behaviors. For example, the experiment performed by Fermi, Pasta, Ulam, and Tsingou reported in [5] produced nonintuitive results and now motivates the Fermi-Pasta-Ulam-Tsingou paradox. We aim to design controls to stabilize these physically relevant fixed grid size systems by driving them towards flocks (see Definition 1).
Although numerous collective behavior models have been introduced and investigated to analyze the emergent flocking of self-organized systems, starting with the pioneering work of Cucker–Smale [2, 3, 8], to the best of our knowledge, this is the first time in the literature that the problem of control around flocks for discrete partial differential equations has been addressed and studied. In recent years, there have been important advances in the control theory for kinetic models, particularly the control to flocking of the Cucker–Smale model [1, 10], which serves as our source of inspiration for studying the problem of controlling the propagation of lattice/chains nonlinear waves.
@@
Controlling discrete nonlinear wave equations for numerical purposes has been well studied in the literature [4, 15]. However, the goal of this research is to design a control that is a good approximation for the continuous problem. Thus, the goal is to observe the behavior of the control as the mesh size tends to zero. Since our problem is meant to be applied to lattices that have a fixed size, there are large differences in the theory. The stochastic control problem for discrete linear wave equations with a fixed mesh size was recently studied in [6, 7]. However, to our knowledge, this work will be the first in the control theory of nonlinear partial differential equations to discuss the fixed mesh size approach for a discrete nonlinear wave equation.
@@ 1.2. Settings and main goals
We consider the following discrete semilinear Klein–Gordon equation with a cubic nonlinearity and a control term
@@
Note that both the semilinear Klein–Gordon equation with cubic and quadratic nonlinearities can be used to describe atomic vibrations in anharmonic lattices [12]. In this work, we only focus on the Klein–Gordon equation with a cubic nonlinearity, while the other case can be treated similarly.
Here, $\alpha,\beta>0$ are given constants. And $w_{0},w_{1}$ are the given initial data, which are $\mathbb{Z}^{n}$-periodic. For $l\in\Lambda$, we write $l^{\prime}\sim l$ if $l^{\prime}-l=\pm he_{k}$ for some $k\in\{1,\ldots,n\}$. The corresponding discrete Laplacian is
@@
where $\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}$ and $\bar{\varphi}$ satisfies (1.2).
We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a stationary flock if
$\varphi(l,t)=\bar{\varphi}(l)\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R},$
where $\bar{\varphi}$ satisfies (1.2).
By symmetry, we only need to consider the moving flocks with $\bar{v}=\sqrt{\alpha/\beta}$. The main questions of interest of our paper are as follows.
@@
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Can we find a control $u\in\mathcal{A}_{M}$ such that $w=\varphi_{1}$ for all $t\geq T$, where $T>0$ is sufficiently large and $\varphi_{1}$ is a moving flock?
@@ 1.3. Main results
with given initial condition
@@
It is worth noting that our ODE system (1.3) is similar to the one studied in [1]. However, the ODE system in [1] does not come from a discrete nonlinear wave equation.
@@
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. There exists $u\in\mathcal{A}_{M}$ such that the solution to (1.3) satisfies,
$\begin{cases}x_{l}(t)=\bar{x}_{1l}+(t-T)\bar{v},\\ v_{l}(t)=\bar{v}=\sqrt{\alpha/\beta},\end{cases}$
for $l\in\overline{\Lambda}$, $t>T$, and some constant $\bar{x}_{1l}\in\mathbb{R}$. In particular, we have $w=\varphi_{1}$ for all $t\geq T$, where $\varphi_{1}$ is a moving flock with
@@
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{2}+a$ for all $t\geq T$.
As a by-product of Theorems 1.11.2, we have the following result.
@@
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{3}+a$ for all $t\geq T$.
Theorem 1.1 can, in some sense, be regarded as the “wave” counterpart to the “particle” results established in the seminal work [1] on controlling the Cucker-Smale model. In contrast, Theorems 1.2 and 1.3 are characteristic of wave phenomena and do not possess a direct “particle” analog. We note that the assumption (1.4) on $M$ in Theorems 1.11.3 is essentially optimal. See Remark 2.
@@
The paper is organized as follows. In Section 2, we prove Theorem 1.1. The proofs of Theorems 1.21.3 are given in Section 3. Finally, section 4 contains a connection between the flocking problem and a mimimal time problem in the framework of optimal control theory and Hamilton–Jacobi equations.
@@
Our strategy to prove Theorem 1.1 is as follows. We start by using a control $u$ that drives the system toward zero velocity. To do so, we first design $u$ so that the particles’ velocities become sufficiently small after some finite time. Once the system is close to zero velocity, we can find a control in $\mathcal{A}_{M}$ that brings it to exactly zero velocity, a stationary flock, after another finite amount of time. Finally, we apply a third type of control to steer the system from a stationary flock to a moving flock with the desired group velocity $\bar{v}=\sqrt{\alpha/\beta}$. It is important noting that all the controls designed are explicit in terms of $(x_{l}(t))_{l\in\overline{\Lambda}}$ and $(v_{l}(t))_{l\in\overline{\Lambda}}$.
@@
which shows that choosing $u_{l}(t)$ we can control $\dot{V}$. For clarity of our writing, we use the usual way to write dot product $(a,b)=ab$ for $a,b\in\mathbb{R}$.
@@
When $t\in[0,T_{0}]$ for some $T_{0}$ we will define later, we will use the control $u_{l}$ based on $v_{l}$ in the following way:
@@
Of course, $u_{l}$ is Lipschitz in $v_{l}$. For practical purposes, we do not need to turn on the control when $|v_{l}|\geq 2a_{2}$. See Remark 1. For each time $t$, we can split the indices $l$ into 3 sets:
@@
Thus, $V(t)$ is nonincreasing and in particular, $V(t)$ is bounded from above by $V(0)$. Since both terms in $V$ are nonnegative, this means they both must be bounded from above. Additionally, for $l\in\overline{\Lambda}$,
@@
Assume this is not the case. Then there exists a $c_{0}>0$, an index $i$, and a sequence of times $t_{k}\to\infty$ such that
@@
In (2.1) above, we define $u_{l}=u_{l}(v_{l})$ in such a way that we do not need to use the control $u_{l}$ when $|v_{l}|\geq 2a_{2}$. This choice is not necessary, but in practical purposes having the control set to zero might be preferred and letting the system be driven by the large velocities in the system might be preferred. Another way, which is simpler but requires a higher magnitude of control, is to define $u_{l}$ as
@@
Since $|\dot{x}_{l}(t)|=|v_{l}(t)|\leq\sqrt{2V(0)}$, the functions $x_{l}(t)$ are Lipschitz with constant $L:=\sqrt{2V(0)}$. Then $x_{l^{\prime}}(t)-x_{l}(t)$ are Lipschitz with Lipschitz constants less than or equal to $2L$. Assume now for the sake of contradiction that (2.7) does not hold. In light of (2.6), there exists an $l$ and a sequence $\{t_{k}\}$ such that $t_{k}\to\infty$ and
@@
=\,= $\left|\left[\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))-((x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))+(x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))}{h^{2}}\right]-c\right|$
[…]
@@
$\begin{cases}|v_{l}(t)|<\eta_{k}<a_{1},\\ |(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)|<\eta_{k}<a_{1}\end{cases}$
@@
$\begin{split}&|\dot{v}_{l}(t_{k}+\tau)-c|\\ =\,&\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t_{k}+\tau)-x_{l}(t_{k}+\tau)}{h^{2}}-c+(\alpha-\beta|v_{l}(t_{k}+\tau)|^{2})v_{l}(t_{k}+\tau)+u_{l}(t_{k}+\tau)\right|\\ \leq\,&8nL\tau+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}.\end{split}$
Using the above inequality, we obtain
@@
[…]
@@
$|v_{l}(t_{k}+\tau)-c\tau|$
[…]
@@
$|c\tau|-|v_{l}(t_{k}+\tau)|\leq\eta_{k}+8nL\tau^{2}+\frac{|c|\tau}{8}+\left(1+\frac{M}{a_{1}}\right)\tau\eta_{k}.$
Now if $\tau=\min\Big{\{}\frac{|c|}{16nL},1\Big{\}}$, then $8nL\tau^{2}\leq\frac{|c|}{2}\tau$, so
@@
However, if we take the limit as $k\to\infty$, we have $\frac{|c|\tau}{4}\leq 0$. Therefore, $c=0$, which is contradicts (2.8). The proof is complete. ∎
@@
It is important noting that $\epsilon$ can be chosen explicit here with the above bound, which is useful in practice. We are now ready to finish the proof of Theorem 1.1.
@@
[…]
=\,= $\epsilon\left(2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta\epsilon^{2}\right)<M.$
@@
[…]
@@
[…]
@@
[…]
@@
$-M\leq-\Delta\bar{\varphi}_{2}(l)<M\quad\text{ for all }l\in\overline{\Lambda}$ (3.4)
@@
$-M<-\Delta\bar{\varphi}_{2}(l)\leq M\quad\text{ for all }l\in\overline{\Lambda}.$ (3.5)
@@
$\begin{split}u_{l}(t)=&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\\ &+\frac{12\alpha}{(T_{4}-T_{3})^{3}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))(t-T_{3})(T_{4}-t)\\ &+\frac{108\sqrt{\alpha\beta}}{(T_{4}-T_{3})^{6}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{2}(t-T_{3})^{2}(T_{4}-t)^{2}\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ &+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}(T_{4}+T_{3}-2t)\\ \geq&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\\ &+\frac{6(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))}{(T_{4}-T_{3})^{3}}\left[(T_{4}+T_{3}-2t)+2\alpha(t-T_{3})(T_{4}-t)\right]\\ &+\frac{216\beta}{(T_{4}-T_{3})^{9}}(\bar{z}+\bar{\varphi}_{2}(l)-x_{l}(T_{3}))^{3}(t-T_{3})^{3}(T_{4}-t)^{3}\\ \geq&\left[-\frac{3(t-T_{3})^{2}}{(T_{4}-T_{3})^{2}}+\frac{2(t-T_{3})^{3}}{(T_{4}-T_{3})^{3}}\right]\Delta\bar{\varphi}_{2}(l)\geq-M.\\ \end{split}$

### v3

# Controlling Klein-Gordon chains and lattices

@@ Abstract.
In this work, we initiate the study of controlling nonlinear Klein–Gordon chains and lattices through their emergent collective flocking behavior. By constructing appropriate feedback control mechanisms, we demonstrate that any physically admissible flock state can be achieved in finite time, meaning the chain can be driven from arbitrary initial vibrations toward a coherent traveling-wave motion. Finally, we reveal a deep connection between the flocking problem and a minimal-time control principle formulated within the framework of nonlinear Hamilton–Jacobi equations and optimal control theory, providing a unifying viewpoint for wave control in discrete nonlinear media.
@@
The work of $S$ is partially supported by NSF grant DMS-2037851. The work of HVT is partially supported by NSF grant DMS-2348305. The work of MBT is partially supported by NSF Grants CAREER DMS-2303146, DMS-2204795, DMS-2306379.
@@ 1.1. Brief overview
Fermi, Pasta, Ulam, and Tsingou [18] investigated how systems evolve toward thermal equilibrium by studying a model consisting of a chain of particles coupled through anharmonic interactions. Their numerical experiment, reported in their original study, produced highly non-intuitive results and eventually led to what is now known as the Fermi–Pasta–Ulam–Tsingou (FPUT) paradox. Although the problem turned out to be far more subtle than they initially expected, their work generated several influential by-products, including the discovery of solitons and breathers, and the realization that FPUT dynamics can, in certain regimes, be approximated by integrable systems. For modern perspectives on these developments, see [1, 2]. The FPUT model is one of the earliest and most renowned examples of an anharmonic lattice. Thus, the FPUT model stands as a classic and foundational system in the study of anharmonic chains and lattices. Progress on the non-equilibrium statistical mechanics of anharmonic chains and lattices remained slow until the late 1990s, when numerical experiments revealed unmistakable signs of anomalous energy transport [3]. Some elements of this behavior had been observed earlier, but only after these findings did extensive, systematic studies begin across many research groups, culminating in the comprehensive reviews [4, 5, 2].
Control theory, developed extensively over the past six decades, has produced deep insights and powerful tools across a wide range of disciplines, including finance, medicine, mechanics, engineering, and climatology [33]. Atomic vibrations in anharmonic chains (one-dimensional systems) and lattices (multi-dimensional systems) are described by discrete wave-type equations [18, 22, 25, 26]. The control of anharmonic chains has emerged as an important research frontier, driving progress in both fundamental science and technological applications (see, for example, [6, 7, 8, 9, 10, 12, 13, 28, 29, 30, 31] and the references therein). This line of research is motivated by the fact that controlling such systems enables the manipulation, redirection, and stabilization of energy flow in discrete media, which are central challenges in condensed matter physics and materials science. Moreover, anharmonic chains are highly nonlinear and often exhibit complex behaviors such as chaos, metastability, and long-time fluctuations. Control mechanisms provide a means to prevent instabilities, suppress unwanted oscillations, stabilize special configurations (including flock states, breathers, and traveling waves), and ensure desirable long-time dynamics. These capabilities are essential for the design of stable devices involving micro- and nanoscale vibrations. Modern engineered systems, including photonic crystals, mechanical metamaterials, and micromechanical resonator arrays, often mirror the structure of anharmonic chains. As a result, control-theoretic tools play a key role in shaping wave propagation in metamaterials, designing energy-harvesting systems, creating tunable filters and waveguides, and managing thermal or mechanical stresses at small scales. Control theory for anharmonic lattices and chains continues to directly influence the development of emerging technologies.
Despite the aforementioned importance in science and technology, rigorous mathematical control designs for anharmonic chains and lattices through the associated discrete wave-type equations have appeared only recently, in [19, 20], where impulsive, feedback, and internal control strategies for harmonic chains were established. Continuing recent studies on the control of anharmonic lattices [19, 20], which focus on linear anharmonic lattices, our work aims to initiate a systematic investigation of how atomic vibrations in nonlinear anharmonic chains and lattices can be controlled through the corresponding nonlinear discrete Klein–Gordon wave equations (Klein–Gordon chains and lattices), following the formulation in [24] (see also [2, 23]).
Because wave vibrations on lattices often display intricate and nonlinear dynamics [11, 27], we focus on designing control mechanisms capable of stabilizing these physically relevant fixed-grid systems by steering them toward flock states, coherent traveling-wave configurations in which particles move collectively with a common velocity direction (see Definition 1). Our work provides the first mathematical foundations for rigorous control designs for nonlinear anharmonic chains and lattices, an important research direction that, despite its significance as discussed above, remains largely unexplored.
@@
We remark that controlling discrete nonlinear wave equations for numerical purposes has been well studied in the literature [17, 34]. However, the goal of this research was to design a control that is a good approximation for the continuous problem. Thus, the goal is to observe the behavior of the control as the mesh size tends to zero. Since our problem is meant to be applied to physical lattices that have a fixed size, there are large differences in the theory.
@@ 1.2. Settings and main goals
We consider the following discrete semilinear Klein–Gordon equation with a cubic nonlinearity following the formulation in [24, Section 2, Equation (2.11) and Section 8, Equation (8.3)] and a control term
@@
Note that the semilinear Klein–Gordon equation with both cubic and quadratic nonlinearities can be used to describe atomic vibrations in anharmonic lattices [24]. In this work, we focus solely on the Klein–Gordon equation with a cubic nonlinearity, while the other case can be treated similarly.
Here, $\alpha,\beta>0$ are given constants, and $w_{0},w_{1}$ are the given initial data, which are $\mathbb{Z}^{n}$-periodic. Let $e_{k}$ for $k\in\{1,\ldots,n\}$ be the standard basis vectors. For $l\in\Lambda$, we write $l^{\prime}\sim l$ if $l^{\prime}-l=\pm he_{k}$ for some $k\in\{1,\ldots,n\}$. The corresponding discrete Laplacian is
@@
where $\bar{v}=\pm\sqrt{\frac{\alpha}{\beta}}$. We say that $\varphi:\overline{\Lambda}\times\mathbb{R}\to\mathbb{R}$ is a stationary flock if
$\varphi(l,t)=\bar{\varphi}(l)\quad\text{ for }(l,t)\in\overline{\Lambda}\times\mathbb{R}.$
By symmetry, we only need to consider the moving flocks with $\bar{v}=\sqrt{\alpha/\beta}$. The main questions of interest in our paper are as follows.
@@
    Let $\bar{v}=\sqrt{\alpha/\beta}$ and $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Can we find a control $u\in\mathcal{A}_{M}$ and an associated $\varphi_{1}$ such that $w=\varphi_{1}$ for all $t\geq T$, where $T>0$ is sufficiently large and $\varphi_{1}$ is a moving flock?
@@ 1.3. Main results
with the given initial condition
@@
Assume (1.4). Let $w_{0},w_{1}$ be $\mathbb{Z}^{n}$-periodic initial data. Then, there exist $u\in\mathcal{A}_{M}$ and a moving flock
@@
where $\bar{x}_{1l}\in\mathbb{R}$ are constants such that $w=\varphi_{1}$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
@@
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{2}+a$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
As a by-product of the proofs of Theorems 1.11.2, we have the following result.
@@
for $l\in\overline{\Lambda}$ and $t>T$. In particular, we have $w=\varphi_{3}+a$ for all $t\geq T$ where $T$ is taken to be sufficiently large.
We end the introduction by noting that in recent years, there have been important advances in control theory for kinetic models, particularly the control to flocking of the Cucker–Smale model [14], which serves as our source of inspiration for studying the problem of controlling the propagation of lattice/chains nonlinear waves. We note that the assumption (1.4) on $M$ in Theorems 1.11.3 is essentially optimal: see Remark 2.
@@
The paper is organized as follows. In Section 2, we prove Theorem 1.1. The proofs of Theorems 1.21.3 are given in Section 3. Finally, section 4 contains a connection between the flocking problem and a minimal time problem in the framework of optimal control theory and Hamilton–Jacobi equations.
@@
Our strategy to prove Theorem 1.1 is as follows. We start by using a control $u$ that drives the system toward zero velocity. To do so, we first design $u$ so that the particles’ velocities become sufficiently small after some finite time. Once the system is close to zero velocity, we can find a control in $\mathcal{A}_{M}$ that brings it to exactly zero velocity, a stationary flock, after another finite amount of time. Finally, we apply a third type of control to steer the system from a stationary flock to a moving flock with the desired group velocity $\bar{v}=\sqrt{\alpha/\beta}$. It is important to note that all the controls designed are explicit in terms of $(x_{l}(t))_{l\in\overline{\Lambda}}$ and $(v_{l}(t))_{l\in\overline{\Lambda}}$.
@@
which shows that by choosing $u_{l}(t)$, we can control $\dot{V}$. For clarity in our writing, we use the usual way to write dot product $(a,b)=ab$ for $a,b\in\mathbb{R}$.
@@
When $t\in[0,T_{0}]$ for some $T_{0}$ to be defined later, we will use the control $u_{l}$ based on $v_{l}$ in the following way:
@@
Of course, $u_{l}$ is Lipschitz in $v_{l}$. See Remark 1 for a discussion on this choice of control. For each time $t$, we can split the indices $l$ into 3 sets:
@@
Thus, $V(t)$ is nonincreasing and, in particular, $V(t)$ is bounded from above by $V(0)$. Since both terms in $V$ are nonnegative, this means they must both be bounded from above. Additionally, for $l\in\overline{\Lambda}$,
@@
Assume this is not the case. Then there exists a $c_{0}>0$, an index $l$, and a sequence of times $t_{k}\to\infty$ such that
@@
In (2.1) above, we define $u_{l}=u_{l}(v_{l})$ in such a way that we do not need to use the control $u_{l}$ when $|v_{l}|\geq 2a_{2}$. This choice is not necessary, but for practical purposes, having the control set to zero might be preferred, and allowing the system to be driven by the large velocities in the system might be preferable. Another way, which is simpler but requires a higher magnitude of control, is to define $u_{l}$ as
@@
Since $|\dot{x}_{l}(t)|=|v_{l}(t)|\leq\sqrt{2V(0)}$, the functions $x_{l}(t)$ are Lipschitz with constant $L:=\sqrt{2V(0)}$. Then $x_{l^{\prime}}(t)-x_{l}(t)$ are Lipschitz with Lipschitz constants less than or equal to $2L$. Assume for the sake of contradiction that (2.7) does not hold. In light of (2.6), there exists an $l$ and a sequence $\{t_{k}\}$ such that $t_{k}\to\infty$ and
@@
$=\,$ $\left|\left[\sum_{l^{\prime}\sim l}\frac{(x_{l^{\prime}}(t)-x_{l}(t))-((x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))+(x_{l^{\prime}}(t_{k})-x_{l}(t_{k}))}{h^{2}}\right]-c\right|$
$\leq\,$ $\frac{8nL}{h^{2}}|t_{k}-t|+\frac{|c|}{8}.$
@@
$\begin{cases}|v_{l}(t)|<\eta_{k}<a_{1},\\ |(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)|<\eta_{k}<a_{1}\end{cases}$ (2.9)
@@
$\begin{split}&|\dot{v}_{l}(t)-c|=\,\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}-c+(\alpha-\beta|v_{l}(t)|^{2})v_{l}(t)+u_{l}(t)\right|\\ \leq\,&\frac{8nL}{h^{2}}|t_{k}-t|+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}.\end{split}$ (2.10)
Let $\tau=\min\{1,\frac{|c|h^{2}}{16nL}\}$. Using (LABEL:31), we obtain
@@
$=\,$ $\left|\int^{t_{k}+\tau}_{t_{k}}\left(\dot{v}_{l}(s)-c\right)\,ds\right|\leq\,\int^{t_{k}+\tau}_{t_{k}}|\dot{v}_{l}(s)-c|\,ds$
$\leq$ $\tau\left(\frac{8nL\tau}{h^{2}}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right)\leq\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right).$
@@
$|v_{l}(t_{k}+\tau)-c\tau|\leq\,|v_{l}(t_{k})|+\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right)$
$\leq\,$ $\eta_{k}+\tau\left(\frac{|c|}{2}+\frac{|c|}{8}+\eta_{k}+\frac{M}{a_{1}}\eta_{k}\right).$
@@
$|c\tau|-|v_{l}(t_{k}+\tau)|\leq\eta_{k}+\frac{|c|\tau}{2}+\frac{|c|\tau}{8}+\left(1+\frac{M}{a_{1}}\right)\tau\eta_{k},$
so
@@
However, if we take the limit as $k\to\infty$, we have $\frac{|c|\tau}{4}\leq 0$. Therefore, $c=0$, which contradicts (2.8). The proof is complete. ∎
@@
It is important to note that $\epsilon$ can be chosen explicitly here with the above bound, which is useful in practice. We are now ready to finish the proof of Theorem 1.1.
@@
$\leq\,$ $\epsilon+\left|\sum_{l^{\prime}\sim l}\frac{x_{l^{\prime}}(t)-x_{l}(t)}{h^{2}}\right|+\alpha\left|v_{l}(T_{0})\right|+\alpha\left|\epsilon(T_{1,l}-T_{0})\right|+\beta\left(\left|v_{l}(T_{0})\right|+\epsilon|T_{1,l}-T_{0}|\right)^{3}$
$\leq\,$ $\epsilon+\frac{8n\epsilon}{h^{2}}+2n\epsilon+\alpha\epsilon+\alpha\epsilon+\beta\left(\left|v_{l}(T_{0})\right|+\epsilon\right)^{3}$
$=\,$ $\epsilon\left(2n+1+\frac{8n}{h^{2}}+2\alpha+8\beta\epsilon^{2}\right)<M.$
@@
$\leq\,$ $\left|\sum_{l^{\prime}\sim l}\frac{\bar{\varphi}_{1}(l^{\prime})-\bar{\varphi}_{1}(l)}{h^{2}}\right|+\frac{12\alpha}{(T_{3}-T_{2})^{3}}|\bar{y}-\bar{\varphi}_{1}(l)|\frac{(T_{3}-T_{2})^{2}}{4}$
@@
$\leq\,$ $\epsilon+\frac{3\alpha|\bar{y}-\bar{\varphi}_{1}(l)|(T_{3}-T_{2})^{2}+6.75\sqrt{\alpha\beta}|\bar{y}-\bar{\varphi}_{1}(l)|^{2}(T_{3}-T_{2})}{(T_{3}-T_{2})^{3}}$
@@
$\leq\,$ $|\Delta\bar{\varphi}_{2}(l)|+\frac{3\alpha|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|(T_{4}-T_{3})^{2}+6.75\sqrt{\alpha\beta}|\bar{\varphi}_{2}(l)-x_{l}(T_{3})|^{2}(T_{4}-T_{3})}{(T_{4}-T_{3})^{3}}$
@@
$M\leq-\Delta\bar{\varphi}_{2}(l)<M\quad\text{ for all }l\in\overline{\Lambda}$ (3.4)
@@
$M<-\Delta\bar{\varphi}_{2}(l)\leq M\quad\text{ for all }l\in\overline{\Lambda}.$ (3.5)
@@
[…]